# Superspin Renormalization and Slow Relaxation in Random Spin Systems

### v1

@@ C.1 Exact Diagonalization
Throughout this work, we perform numerical ED of the system Hamiltonians (1, (40)) at even number of sites. Crucially, we reduce the Hilbert space dimension by the non-interacting nature for the nearest-neighbor models (Appendix C.1.1) and by the $\mathrm{U}(1)$ and $\mathbb{Z}_{2}$ symmetries (2) for the beyond-nearest-neighbor and two-dimensional models (Appendix C.1.2). With the exactly diagonalized Hamiltonian, we compute the $\overline{\mathcal{S}_{p}}(\omega)$ of interest at low frequencies. For numerical convenience, in our calculations of $\overline{S_{p}}(t)$ in Fig. 5, we neglect contributions from pairs of eigenstates with energy separations greater than a UV cutoff $\epsilon_{\mathrm{UV}}=1$. While this leads to errors at early times, it does not significantly affect the late-time behavior that we are interested in.
@@ C.2 Numerical RSRG-X
In principle, given a realization of the randomness we can numerically track $H_{\mathrm{eff}}$ along all the possible branching choices detailed for the RG steps depending on the specific models (Sec. III.2, Sec. V.1). In this way, we can obtain the complete set of RSRG-X states with the corresponding approximate eigenenergies. The information so obtained would then enable us to simulate the late-time dynamics, spin relaxation for example. However, this straightforward approach is computationally inefficient as the total number of RSRG-X states grows exponentially with system sizes. We simplify our approach to various extents respectively for the nearest-neighbor models in Appendix C.2.1 and for the general interacting models — including specifically chains with interactions beyond the nearest neighbors (1) and the two-dimensional system (40) — in Appendix C.2.2.
@@ C.2.2 Beyond-Nearest-Neighbor and Two-Dimensional Models
Within the assumption of strong randomness, we have argued that the autocorrelation function can be approximated using Eqs. (38) and (39). The approximation leads to a spectral function
@@ D.1 General Perturbative Treatment
We outline our general strategy before detailing the perturbative treatment for every choice of branching subspace $\mathcal{V}$. Our goal is to describe the updated Hamiltonian $H_{\mathrm{eff}}^{(\mathcal{V})}$ due to branching into $\mathcal{V}$, given the $H_{\mathrm{eff}}$ an RG step begins with: as in Equ. (8), we separate $H_{\mathrm{eff}}$ into the strong interaction $H_{0}$, the interactions $H_{1}$ (other than $H_{0}$) that overlap with $H_{0}$ at some site, and the interactions $H_{\perp}$ that do not involve any sites supporting $H_{0}$. For convenience, let $H_{0}$ be supported on superspin site(s) indexed by an index set $\mathbf{S}$ (e.g., $\mathbf{S}=\{1,2\}$ in cases (i), (ii), and $\mathbf{S}=\{1\}$ in case (iii) in Sec. V.1.3). It follows that $H_{\perp}$ is supported on $\overline{\mathbf{S}}$, by which we denote all the other superspin sites. Eventually, we express $H_{\mathrm{eff}}^{(\mathcal{V})}$ in separate terms as are contributed by the perturbation theory (9),

### v2

@@ C.1 Exact Diagonalization
Throughout this work, we perform numerical ED of the system Hamiltonians (140) at even number of sites. Crucially, we reduce the Hilbert space dimension by the non-interacting nature for the nearest-neighbor models (Appendix C.1.1) and by the $\mathrm{U}(1)$ and $\mathbb{Z}_{2}$ symmetries (2) for the beyond-nearest-neighbor and two-dimensional models (Appendix C.1.2). With the exactly diagonalized Hamiltonian, we compute the $\overline{\mathcal{S}_{p}}(\omega)$ of interest at low frequencies. For numerical convenience, in our calculations of $\overline{S_{p}}(t)$ in Fig. 5, we neglect contributions from pairs of eigenstates with energy separations greater than a UV cutoff $\epsilon_{\mathrm{UV}}=1$. While this leads to errors at early times, it does not significantly affect the late-time behavior that we are interested in.
@@ C.2 Numerical RSRG-X
In principle, given a realization of the randomness we can numerically track $H_{\mathrm{eff}}$ along all the possible branching choices detailed for the RG steps depending on the specific models (Secs. III.2V.1). In this way, we can obtain the complete set of RSRG-X states with the corresponding approximate eigenenergies. The information so obtained would then enable us to simulate the late-time dynamics, spin relaxation for example. However, this straightforward approach is computationally inefficient as the total number of RSRG-X states grows exponentially with system sizes. We simplify our approach to various extents respectively for the nearest-neighbor models in Appendix C.2.1 and for the general interacting models — including specifically chains with interactions beyond the nearest neighbors (1) and the two-dimensional system (40) — in Appendix C.2.2.
@@ C.2.2 Beyond-Nearest-Neighbor and Two-Dimensional Models
Within the assumption of strong randomness, we have argued that the autocorrelation function can be approximated using Eqs. (3839). The approximation leads to a spectral function
@@ D.1 General Perturbative Treatment
We outline our general strategy before detailing the perturbative treatment for every choice of branching subspace $\mathcal{V}$. Our goal is to describe the updated Hamiltonian $H_{\mathrm{eff}}^{(\mathcal{V})}$ due to branching into $\mathcal{V}$, given the $H_{\mathrm{eff}}$ an RG step begins with: as in Equ. (8), we separate $H_{\mathrm{eff}}$ into the strong interaction $H_{0}$, the interactions $H_{1}$ (other than $H_{0}$) that overlap with $H_{0}$ at some site, and the interactions $H_{\perp}$ that do not involve any sites supporting $H_{0}$. For convenience, let $H_{0}$ be supported on superspin site(s) indexed by an index set $\mathbf{S}$ [e.g., $\mathbf{S}=\{1,2\}$ in cases (i), (ii), and $\mathbf{S}=\{1\}$ in case (iii) in Sec. V.1.3]. It follows that $H_{\perp}$ is supported on $\overline{\mathbf{S}}$, by which we denote all the other superspin sites. Eventually, we express $H_{\mathrm{eff}}^{(\mathcal{V})}$ in separate terms as are contributed by the perturbation theory (9),